# A Heavy Seed Black Hole Mass Function at High Redshift - Prospects for LISA

### v1

@@ 1 Introduction
To this end models have focused on trying to explain how MBHs can form and ultimately end up powering the quasars in galactic nuclei. There are a number of potential pathways that can lead to a MBH (e.g. Rees, 1978). In broad terms the pathways can be broken into the light seed pathway and the heavy seed pathway. The light seed pathway (e.g. Madau & Rees, 2001) relies on the remnants of Population III (Pop III) stars, the first generations of stars in the Universe. The black holes formed from these stars would have an upper limit of at most 103 M (Schneider et al., 2002; Hartwig et al., 2015; Lazar & Bromm, 2022). In order for these black holes to reach the masses that we observe at high-z, they would need to undergo accretion at the Eddington rate for extended periods and for the high-z quasar population, in particular, accretion at, or beyond, the Eddington rate with a duty cycle of very close to unity. This may be physically unlikely as the black hole would need a consistent supply of matter to accrete from and would need to stay at the centre of a dense accretion flow (Whalen, Abel & Norman, 2004; Milosavljević, Couch & Bromm, 2009; Alvarez, Wise & Abel, 2009). The dynamics of black hole accretion has turned out to be one of the main bottlenecks to efficient accretion onto light seed black holes with numerous studies showing that light seed black holes are unable to efficiently sink to the high-density centre of galaxies (e.g. Pfister et al., 2019; Ma et al., 2021). While dynamics, without doubt, does provide a strong bottleneck to efficient growth, growth within more favourable environment may be possible (e.g. Lupi et al., 2014; Alexander & Natarajan, 2014; Lupi et al., 2016; Natarajan, 2020; Shi et al., 2023; Shi, Kremer & Hopkins, 2024, 2024; Gordon et al., 2024; Mehta, Regan & Prole, 2024). The question is how frequently this efficient (possibly via bursts of super/hyper-Eddington) accretion occurs and what is the parameter space that allows it to occur? If the probability of growth is too low then perhaps another seed formation pathway is required.
@@
The heavy seed pathway emerges as an attractive scenario in this case. For the heavy seed pathway the initial mass of the black holes is expected to be in the range $10^{3}-10^{5}\ \rm{M}_{\odot}$. The actual formation of the heavy seed can occur via two, potentially overlapping, pathways. On the one hand we have formation via a supermassive star (SMS) in the nucleus of the halo (e.g. Latif et al., 2013a, b; Latif, Schleicher & Hartwig, 2016; Regan & Downes, 2018; Woods et al., 2019; Regan et al., 2020), or else a heavy seed may form via dynamical runaway within a dense stellar cluster or dense black hole cluster (Devecchi & Volonteri, 2009; Katz, Sijacki & Haehnelt, 2015; Rizzuto et al., 2021; González et al., 2021; Fragione & Leigh, 2018; Fragione et al., 2022; Antonini, Gieles & Gualandris, 2019; Mapelli et al., 2021; Arca Sedda, Amaro Seoane & Chen, 2021; Reinoso et al., 2023). Both of these scenarios have been thoroughly investigated via analytic, semi-analytic and numerical modeling over the last decade. While convergence of models is certainly lacking it is also true that heavy seed formation has been shown to be viable and with number density values shown to be potentially compatible with the observed population of high-z MBHs (Bogdán et al., 2023).
@@ 2.1 The Renaissance Simulations
The goal of the Renaissance simulation suite was to model early galaxy formation, including Pop III and Pop II star formation and feedback. In order to accomplish this, high resolution was required to allow us to resolve (mini-)halos of masses $\sim 10^{6}\ \rm{M}_{\odot}$ and then track their growth through accretion and mergers with other halos. The adaptive mesh refinement abilities that Enzo possesses allows for the increase of resolution for select, dense regions within the simulation box. Due to the computational expense of such a strategy it could not be employed over the entire parent box (40 cMpc). Instead three regions were chosen within the parent box to represent different regions that host galaxy formation of different large-scale overdensity. The three regions are the Rarepeak (RP) (run to a redshift of $z=15$), Normal (run to $z=11.6$) and Void (run to $z=8$) regions (Chen et al., 2014), the names refering to their level of overdensity with respect to the cosmic mean. As the goal of this study is to constrain the number density of heavy seed black holes we choose to study only the Normal and Rarepeak regions here as the Void region has been previously found to support only very few, if any, heavy seeds (Regan et al., 2020).
The Renaissance simulations contain models for metal free (Population III) and metal-poor (Population II) to allow for self consistent (early) star formation. Crucially the simulations, however, do not contain a prescription for black hole formation (see also Smith et al., 2018a). We will use the information that the Renaissance simulation snapshots supply to predict regions in which massive black hole (MBH) formation is likely to occur. We then use that same information to predict the accretion and merger rates for MBHs in the early Universe.
@@ 2.2 Seeding the Massive Black Hole Population
Finally, we consider only halos with a mass inflow rate exceeding 0.1 $\rm{M_{\odot}\ yr^{-1}}$ which is needed to generate the conditions thought necessary to allow for (super-)massive stars to form (Hosokawa, Omukai & Yorke, 2012; Woods et al., 2019; Wise et al., 2019; Regan, 2023), thus leading into a heavy seed. The mass inflow rate is directly connected to the growth rate of the halo (see e.g. Regan, 2023) with higher growth rates leading to higher dynamical heating within the halo (Yoshida et al., 2003)
@@
where $\sigma$ is the sphere of virial radius $R_{vir}$ centered on the highest density point of the halo in question (Bleuler & Teyssier, 2014; Regan et al., 2020). The inflow is calculated on a sphere of 20 pc radius centered on the highest density point. We choose the highest density point as the centre due to the clumpy nature of the halo.
Once all of the heavy seed halo conditions are met, we assume the formation of a heavy seed MBH. In reality the heavy seed MBH will be preceded by phase of stellar evolution either via a supermassive star, a dense stellar cluster or some combination. In either case the lifetime of the stellar phase of the heavy seed is expected to be of order $10^{6}$ years (Yungelson et al., 2008; Woods et al., 2019). Due to the fact that the time between snapshots in Renaissance is on the order of a few million years, we can assume that our heavy seeds can immediately be identified as MBHs.
@@ 2.3 Assigning Masses to the Seeds
For the Pacucci et al. (2023) relation, the free parameters are set to be $\gamma=1.9$ and $\delta=-11.9$. While for the Reines & Volonteri (2015) relation, the free parameters are set to $\gamma=1.9$ and $\delta=-13.78$ respectively. We hereafter refer to the seeding method using the relation from Reines & Volonteri (2015) as RV15 and to the method using Pacucci et al. (2023) as P23. Finally, we impose a minimum black hole seeding mass of MBH > 100 M. We include black holes with masses less than $10^{3}\rm{M}_{\odot}$ to investigate whether light seeds in lower mass galaxies are able to grow enough in order to reach masses comparable to heavy seed black holes.
@@ 2.4 Growth of the MBH population
The next question is where to position our seeded MBHs at each timestep (i.e. at every output dataset). For this we take two different approaches. In the first case we simply place the MBH at the point of maximum (gas) density in each host halo, the $\rho_{max}$ method. Taking the highest density point as the default position of the MBH is clearly a somewhat optimistic approach. Due to the relatively low mass of even the most massive galaxies in Renaissance (${}_{*,max}\sim 5\times 10^{7}\ \rm{M}_{\odot}$) the central nucleus of the halo will be poorly defined and the central potential relatively shallow. Moreover, the mass of the seed black hole (${}_{\rm{seed}}\lesssim 10^{4}\text{ M}_{\odot}$) does not guarantee that it will sink to the center of the halo. In fact, it has been shown from simulations that MBHs with masses below $10^{4}\text{M}_{\odot}$ do not sink to the center within a Hubble time (e.g. Pfister et al., 2019; Ma et al., 2021). This is therefore the most optimistic and simplest approach and almost certainly not wholly realistic. To counter this we also take a second approach.
@@ 2.5 Boosting the Accretion
As can be seen from eq.(8), we include a boost factor, $\alpha$, to account for the fact that Renaissance may underestimate the accretion onto a MBH, in unresolved regions where the cell length is larger than the Bondi radius described by eq.(9). The population of black holes where the Bondi radius is unresolved/resolved is seen in Figure 2.
Figure 2: The population of Massive Black Holes showing the ratio between the Bondi radius, $r_{b}$, and the cell length, $\Delta x$, for the two black hole positions, for each region from Renaissance. The masses for the black holes are from P23. Using the RV15 method would only make the Bondi radii less resolved, thus further requring the boost factor for the accretion calculation. The z-axis represents the number density of gas in the region surrounding the black hole. The grey region represents the black holes where the Bondi radius is unresolved.
@@ 2.7 MBH Merger Delay Times
We also assume that the stellar mass of the halo remains constant as the black holes move through it. In reality and depending on the time scale of the MBH binary merger this may not be the case. Nonetheless, the approximation gives us a sense of time scale for the binary merger and particulatly whether or not we would expect the binary to merge prior to z = 0.
@@ 3.1 MBH Host Halo Identification
Using the metadata from the Normal and RP regions from Renaissance, we are able to search for halos which satisfy the metallicity conditions, inflow rate conditions and compactness conditions. In Figure 3 we show the number of halos that satisfy each and all of these conditions. The top panel of Figure 3 is for the RP region, the bottom panel for the Normal region. In each case only haloes with masses in excess of the atomic cooling limit are considered (condition 1, see §2.2). Following this, the metallicity condition (Z $<10^{-3}\rm{\ Z_{\odot}}$) is the easiest condition to satisfy (green line). The next “easiest” condition to satisfy is the compactness condition followed by the inflow rate condition. The black line in each panel gives the total number of haloes satisfying our conditions for heavy seed production at each timestep.
Once a halo is seeded with a black hole, it cannot be seeded again. From Figure 3 it is clear that for overdense regions such as the RP region the number of heavy seeds forming is greater overall and that the haloes are also seeded much earlier. Out of all the halos identified in the RP region using Rockstar, 371 ($\sim$ 4 % of atomic cooling halos) are seeded with central MBHs (by z = 15). For the Normal region, 53 ($\sim$ 1%) are seeded with central MBHs (by z = 11.6).
@@ 3.3 The Growth of Heavy Seed MBHs in Renaissance
We examine the accretion onto the MBHs by computing the Eddington fraction, $f_{\text{edd}}=\dot{m}/\dot{m}_{\text{edd}}$, for both regions’ (i.e. Rarepeak & Normal) populations. In Figure 5 we show 2-D hisrograms of accretion (as a fraction of the Eddington rate) as a function of black hole mass. The accretion rate shown is the maximum rate across the entire redshift range and hence does not truly reflect the mass growth but instead gives a flavour for the maximum instantaneous growth rates. The top panel shows maximum accretion rate from the RP region while the bottom panels show the maximum instantaneous accretion rate from the Normal region. The left column uses MBHs with initial masses assigned from the P23 relationship, the right hand panel shows initial MBH masses assigned from the RV15 relationship. In orange we show the result when the black holes are positioned consistently at the highest density point while in blue we show the results for when the MBHs follow the dynamics due to the initially nearest DM particle.
@@
Figure 6: The SNR of the mergers, for the no-delay scenario, found in the Renaissance simulations, with the SNR of the mergers calculated from Robson, Cornish & Liu (2019). The black crosses represent the mergers where the SNR could not be calculated and the black dashed line represents the case where the mass ratio between the two merging black holes is 1. The grey shaded regions denote BH masses $<10^{3}\ \rm{M}_{\odot}$.
Figure 7: The merger times of the black holes from Renaissance with the effects of dynamical friction taken into account. The grey shaded region denotes mergers that would not be detected due to the fact that they would not have occurred yet. The hatched regions denote BH masses $<10^{3}\ \rm{M}_{\odot}$. The region of interest, or at least the region of detectibility with LISA, is then the bottom right quadrant.

### v2

@@ 1 Introduction
To this end models have focused on trying to explain how MBHs can form and ultimately end up powering the quasars in galactic nuclei. There are a number of potential pathways that can lead to a MBH (e.g. Rees, 1978). In broad terms the pathways can be broken into the light seed pathway and the heavy seed pathway. The light seed pathway (e.g. Madau & Rees, 2001) relies on the remnants of Population III (Pop III) stars, the first generations of stars in the Universe. The black holes formed from these stars would have an upper limit of at most 103 M (Schneider et al., 2002; Hartwig et al., 2015; Lazar & Bromm, 2022). In order for these black holes to reach the masses that we observe at high-z, they would need to undergo accretion at the Eddington rate for extended periods and for the high-z quasar population, in particular, accretion at, or beyond, the Eddington rate with a duty cycle of very close to unity. This may be physically unlikely as the black hole would need a consistent supply of matter to accrete from and would need to stay at the centre of a dense accretion flow (Whalen, Abel & Norman, 2004; Milosavljević, Couch & Bromm, 2009; Alvarez, Wise & Abel, 2009).
The dynamics of black hole accretion has turned out to be one of the main bottlenecks to efficient accretion onto light seed black holes with numerous studies showing that light seed black holes are unable to efficiently sink to the high-density centre of galaxies (e.g. Pfister et al., 2019; Ma et al., 2021). While dynamics, without doubt, does provide a strong bottleneck to efficient growth, growth within more favourable environment may be possible (e.g. Lupi et al., 2014; Alexander & Natarajan, 2014; Lupi et al., 2016; Natarajan, 2020; Shi et al., 2023; Shi, Kremer & Hopkins, 2024, 2024; Gordon et al., 2024; Mehta, Regan & Prole, 2024). The question is how frequently this efficient (possibly via bursts of super/hyper-Eddington) accretion occurs and what is the parameter space that allows it to occur? If the probability of growth is too low then perhaps another seed formation pathway is required.
@@
The heavy seed pathway emerges as an attractive scenario in this case. For the heavy seed pathway the initial mass of the black holes is expected to be in the range $10^{3}-10^{5}\ \rm{M}_{\odot}$. The actual formation of the heavy seed can occur via two, potentially overlapping, pathways. On the one hand we have formation via a supermassive star (SMS) in the nucleus of the halo (e.g. Latif et al., 2013a, b; Latif, Schleicher & Hartwig, 2016; Regan & Downes, 2018; Woods et al., 2019; Regan et al., 2020), or else a heavy seed may form via dynamical runaway within a dense stellar cluster or dense black hole cluster (Devecchi & Volonteri, 2009; Katz, Sijacki & Haehnelt, 2015; Rizzuto et al., 2021; González et al., 2021; Fragione & Leigh, 2018; Fragione et al., 2022; Antonini, Gieles & Gualandris, 2019; Mapelli et al., 2021; Arca Sedda, Amaro Seoane & Chen, 2021; Reinoso et al., 2023). Both of these scenarios have been thoroughly investigated via analytic, semi-analytic and numerical modeling over the last decade. While convergence of models is certainly lacking it is also true that heavy seed formation has been shown to be viable and with number density values shown to be potentially compatible with the observed population of high-z MBHs (Bogdán et al., 2023). More exotic channels for black hole formation include the formation of massive protostars that are fueled by the annihilation of dark matter particles instead of nuclear fusion Spolyar, Freese & Gondolo (2008).
Over the last two decades a number of channels have been explored which can in principle lead to environments capable of supporting heavy seed formation. These are i) formation of heavy seeds in haloes exposed to large Lyman-Werner (LW) fields, ii) baryonic streaming velocities leading to heavy seed formation and iii) rapid halo assembly leading to heavy seed formation. In this paper we focus exclusively on the rapid-assembly channel. Further details on the other channels can be found in the review by Inayoshi, Visbal & Haiman (2020). The rapid-assembly channel potentially leads to a much greater number density of MBHs at high-z being only dependent on the assembly history of the halo up to the atomic cooling limit. Recent numerical simulations by Wise et al. (2019), Regan et al. (2020), Lupi, Haiman & Volonteri (2021) and Latif et al. (2022) shown that inflow rates of approximately 0.1 $\rm{M_{\odot}\ yr^{-1}}$ into the halo centre (i.e. down to scales of order 1 - 50 pc) is sufficient to suppress the cooling effect of $\rm{H_{2}}$ allowing larger Jeans masses to develop. In turn this allows the formation of massive proto-stars (Regan et al., 2020; Prole et al., 2024). It is this pathway we explore in this paper and from this model we quantify the expected number density of heavy seeds.
@@ 2.1 The Renaissance Simulations
The goal of the Renaissance simulation suite was to model early galaxy formation, including Pop III and Pop II star formation and feedback. In order to accomplish this, high resolution was required to allow us to resolve (mini-)halos of masses $\sim 10^{6}\ \rm{M}_{\odot}$ and then track their growth through accretion and mergers with other halos. The adaptive mesh refinement abilities that Enzo possesses allows for the increase of resolution for select, dense regions within the simulation box. Due to the computational expense of such a strategy it could not be employed over the entire parent box (40 cMpc). Instead three regions were chosen within the parent box to represent different regions that host galaxy formation of different large-scale overdensity. The three regions are the Rarepeak (RP) (run to a redshift of $z=15$ with dimensions of 3.8 × 5.4 × 6.6 Mpc3), Normal (run to $z=11.6$, 6.0 × 6.0 × 6.125 Mpc3) and Void (run to $z=8$, with dimensions the same as the Normal region) regions (Chen et al., 2014), the names referring to their level of overdensity with respect to the cosmic mean. As the goal of this study is to constrain the number density of heavy seed black holes we choose to study only the Normal and Rarepeak regions here as the Void region has been previously found to support only very few, if any, heavy seeds (Regan et al., 2020).
The Renaissance simulations contain models for metal free (Population III) and metal-poor (Population II) to allow for self consistent (early) star formation. Crucially the simulations, however, do not contain a prescription for black hole formation (see also Smith et al., 2018a). We will use the information that the Renaissance simulation snapshots supply to predict regions in which massive black hole (MBH) formation is likely to occur. We then use that same information to predict the accretion and merger rates for MBHs in the early Universe. The resolution for the dark matter particles in Renaissance is $2.4\times 10^{4}\rm{M}_{\odot}$ while the gravitational softening length is resolved down as far as 19pc.
@@ 2.2 Seeding the Massive Black Hole Population
Finally, we consider only halos with a mass inflow rate exceeding 0.1 $\rm{M_{\odot}\ yr^{-1}}$ which is needed to generate the conditions thought necessary to allow for (super-)massive stars to form (Hosokawa, Omukai & Yorke, 2012; Woods et al., 2019; Wise et al., 2019; Regan, 2023), thus leading into a heavy seed. The gas mass inflow rate is directly connected to the growth rate of the halo (see e.g. Regan, 2023) with higher growth rates leading to higher dynamical heating within the halo (Yoshida et al., 2003)
@@
where $\sigma$ is the sphere of virial radius $R_{vir}$ centered on the highest density point of the halo in question, $\vec{v}$ is the velocity of the gas flowing through an infinitesimal area element on $\sigma$ (Bleuler & Teyssier, 2014; Regan et al., 2020). The inflow is calculated on a sphere of 20 pc radius centered on the highest density point. We choose the highest density point as the centre due to the clumpy nature of the halo.
Once all of the heavy seed halo conditions are met, we assume the formation of a heavy seed MBH. In reality the heavy seed MBH will be preceded by phase of stellar evolution either via a supermassive star, a dense stellar cluster or some combination. In either case the lifetime of the stellar phase of the heavy seed is expected to be of order $10^{6}$ years (Yungelson et al., 2008; Woods et al., 2019). Due to the fact that the time between snapshots in Renaissance is on the order of a few million years, we can assume that our heavy seeds can immediately be identified as MBHs. It is noted that because we allow heavy seeds to form in halos that reach atomic cooling limit ($\sim 10^{6}\rm{M}_{\odot}$, along with the other criteria for seeding), this allows for black holes to form in earlier times compared to other simulations that seed at $10^{10}\rm{M}_{\odot}$ (EAGLE McAlpine et al. (2016), FABLE Henden et al. (2018) and SIMBA Davé et al. (2019)). Thus, giving the black holes in this model extra time to grow in comparison to other cosmological simulation.
Some of our choices for the seeding prescription are open to investigation (as is almost universally true for all subgrid models). In particular our choice of the Inflow Rate threshold (0.1 $\rm{M_{\odot}\ yr^{-1}}$ ) and the compactness parameter are valid avenues for further exploration. The inflow rate threshold is the parameter that ultimately dictates the number density (see Figure 3). The value chosen is based on the high resolution studies of Regan et al. (2020) and Regan (2023) which show that MBH seeds can form in haloes experiencing inflow rates at or above this threshold. As such we believe our choice to be robust.
@@ 2.3 Assigning Masses to the Seeds
For the Pacucci et al. (2023) relation, the free parameters are set to be $\gamma=1.9$ and $\delta=-11.9$. While for the Reines & Volonteri (2015) relation, the free parameters are set to $\gamma=1.9$ and $\delta=-13.78$ respectively. We hereafter refer to the seeding method using the relation from Reines & Volonteri (2015) as RV15 and to the method using Pacucci et al. (2023) as P23. Finally, we impose a minimum black hole seeding mass of MBH > 100 M. We include black holes with masses less than $10^{3}\rm{M}_{\odot}$ to investigate whether light seeds in lower mass galaxies are able to grow enough in order to reach masses comparable to heavy seed black holes. We decided to use a range of black hole seed masses, in contrast to other simulations (e.g. EAGLE McAlpine et al. (2016), FABLE Henden et al. (2018) and SIMBA Davé et al. (2019)). Using a range of masses will allow us to study the impacts black hole mass will have on the accretion calculations onto the black holes as well as the kind of mass ratios we expect to see in black hole mergers.
@@ 2.4 Growth of the MBH population
The next question is where to position our seeded MBHs at each timestep (i.e. at every output dataset). For this we take two different approaches. In the first case we simply place the MBH at the point of maximum (gas) density in each host halo and keep it there at each timestep, the $\rho_{max}$ method. The highest density point is recorded independently of the black holes existence via the halo merger tree, so there is no chance of the black hole moving to a nearby galaxy that contains a more dense region than the original halo. Taking the highest density point as the default position of the MBH is clearly a somewhat optimistic approach in this model. Realistically, however, this method may not be the most optimistic approach, due to feedback effects from star formation in the dense local environment. This stellar feedback may effectively expel a substantial amount of the gas from the local reservoir, that would otherwise be accreted by the massive black hole still in the dense environment (Dubois et al., 2015). There is also the matter of feedback from the black hole itself during its accretion cycle, which could also impact the surrounding gas density, thus leading to slight inaccuracies in the accretion calculation, see Sijacki et al. (2015); Schaye et al. (2015); Bennett et al. (2023); Ni et al. (2022); Trebitsch et al. (2022) Due to the relatively low mass of even the most massive galaxies in Renaissance (${}_{*,max}\sim 5\times 10^{7}\ \rm{M}_{\odot}$) the central nucleus of the halo will be poorly defined and the central potential relatively shallow. Moreover, the mass of the seed black hole (${}_{\rm{seed}}\lesssim 10^{4}\text{ M}_{\odot}$) does not guarantee that it will sink to the center of the halo. In fact, it has been shown from simulations that MBHs with masses below $10^{4}\text{M}_{\odot}$ do not sink to the center within a Hubble time (e.g. Pfister et al., 2019; Ma et al., 2021). This is therefore the most optimistic and simplest approach and almost certainly not wholly realistic. To counter this we also take a second approach.
@@ 2.5 Boosting the Accretion
As can be seen from eq.(8), we include a boost factor, $\alpha$, to account for the fact that Renaissance may underestimate the accretion onto a MBH, in unresolved regions where the cell length is larger than the Bondi radius described by eq.(9). Only MBHs where the Bondi radius is not resolved are boosted. The population of black holes where the Bondi radius is unresolved/resolved is seen in Figure 2.
Figure 2: The population of Massive Black Holes showing the ratio between the Bondi radius, $r_{b}$, and the cell length, $\Delta x$, for the two black hole positions, for each region from Renaissance. The masses for the black holes are from P23. Using the RV15 method would only make the Bondi radii less resolved, thus further requiring the boost factor for the accretion calculation. The z-axis represents the number density of gas in the region surrounding the black hole. The grey region represents the black holes where the Bondi radius is unresolved.
@@ 2.7 MBH Merger Delay Times
We also assume that the stellar mass of the halo remains constant as the black holes move through it. In reality and depending on the time scale of the MBH binary merger this may not be the case. Nonetheless, the approximation gives us a sense of time scale for the binary merger and particularly whether or not we would expect the binary to merge prior to z = 0.
@@ 3.1 MBH Host Halo Identification
Using the metadata from the Normal and RP regions from Renaissance, we are able to search for halos which satisfy the metallicity conditions, inflow rate conditions and compactness conditions. In Figure 3 we show the number of halos that satisfy each and all of these conditions. The solid lines in Figure 3 are for the RP region, the dashed lines for the Normal region. The strictest condition for the identifying haloes is the mass inflow criteria. This is the criteria which dominates the seeding identification. In each case only haloes with masses in excess of the atomic cooling limit are considered (condition 1, see §2.2). Following this, the metallicity condition (Z $<10^{-3}\rm{\ Z_{\odot}}$) is the easiest condition to satisfy (green line). The next “easiest” condition to satisfy is the compactness condition followed by the inflow rate condition. The black line in each panel gives the total number of haloes satisfying our conditions for heavy seed production at each timestep.
Once a halo is seeded with a black hole, it cannot be seeded again. From Figure 3 it is clear that for overdense regions such as the RP region the number of heavy seeds forming is greater overall and that the haloes are also seeded much earlier at $z\sim 22$, in contrast to the Normal region which starts seeding at $z\sim 16$. Because of the criteria of the halo to reach atomic cooling limit ($\sim 10^{6}\rm{M}_{\odot}$), and the capabilities of Renaissance to resolve halos of this mass, we can be assured that Renaissance is able to capture the possible halos that are seeding black holes at high redshifts. Out of all the halos identified in the RP region using Rockstar, 371 ($\sim$ 4 % of atomic cooling halos) are seeded with central MBHs (by z = 15). For the Normal region, 53 ($\sim$ 1%) are seeded with central MBHs (by z = 11.6). We note that the conditions required for seeding show no signs of slowing down after the cessation of each region, leaving the possibility of additional black holes being formed in these high redshift regimes.
It is clear from Figure 3 that the level of isolation that permits black holes from seeding in Renaissance plays an important role. We see that an overdense region, such as RP, is able to produce black holes earlier and at an order of magnitude higher at its simulation end time compared to that of the Normal region. In contrast, the Void region, which exists at a higher level of isolation fails to produce any halos capable of seeding heavy seed black holes.
@@
Figure 4: Histogram showing the distribution of virial masses of the host halos at the time of their respective black hole seeding.
@@ 3.3 The Growth of Heavy Seed MBHs in Renaissance
We examine the accretion onto the MBHs by computing the Eddington fraction, $f_{\text{edd}}=\dot{m}/\dot{m}_{\text{edd}}$, for both regions’ (i.e. Rarepeak & Normal) populations. In Figure 6 we show 2-D histograms of accretion (as a fraction of the Eddington rate) as a function of black hole mass. The accretion rate shown is the maximum rate across the entire redshift range and hence does not truly reflect the mass growth but instead gives a flavour for the maximum instantaneous growth rates. The top panel shows maximum accretion rate from the RP region while the bottom panels show the maximum instantaneous accretion rate from the Normal region. The left column uses MBHs with initial masses assigned from the P23 relationship, the right hand panel shows initial MBH masses assigned from the RV15 relationship. In orange we show the result when the black holes are positioned consistently at the highest density point while in blue we show the results for when the MBHs follow the dynamics due to the initially nearest DM particle.
@@
Figure 9: The SNR of the mergers, for the no-delay scenario, found in the Renaissance simulations, with the SNR of the mergers calculated from Robson, Cornish & Liu (2019). Due to the lack of growth for the black holes, the $\rho_{max}$ methods and the DMproxy methods are included in the same markers. The black crosses represent the mergers where the SNR could not be calculated and the black dashed line represents the case where the mass ratio between the two merging black holes is 1. The grey shaded regions denote BH masses $<10^{3}\ \rm{M}_{\odot}$.
Figure 10: The merger times of the black holes from Renaissance with the effects of dynamical friction taken into account. The separation is measured in proper units. As in Figure 9, due to the lack of growth for the MBHs, the $\rho_{max}$ methods and the DMproxy methods are included in the same markers. The grey shaded region denotes mergers that would not be detected due to the fact that they would not have occurred yet. The hatched regions denote BH masses $<10^{3}\ \rm{M}_{\odot}$. The region of interest, or at least the region of detectability with LISA, is then the bottom right quadrant.